# Leveraging Large Language Model for Information Retrieval-based Bug Localization

### v1

@@ IX Conclusion and Future Work
This paper presents GenLoc, a novel information retrieval-based bug localization technique that integrates semantic retrieval with step-by-step reasoning capabilities of LLMs. An evaluation on over 9,000 real-world bugs from six diverse Java projects shows that GenLoc outperforms existing techniques in terms of Accuracy@k, MRR@10, and MAP@10. The results highlight GenLoc’s ability to localize bugs that lack direct lexical overlap with code, a major challenge for conventional IR-based methods. Moreover, GenLoc can localize the highest number of unique bugs, demonstrating its potential to complement existing approaches.
In future, GenLoc could be extended to support method-level bug localization to provide finer-grained debugging assistance. Furthermore, the impact of different LLMs on both the effectiveness and efficiency of localization could be analyzed.

### v2

@@ 9. Conclusion and Future Work
This paper presents GenLoc, a novel information retrieval-based bug localization technique that integrates semantic retrieval with step-by-step reasoning capabilities of LLMs. An evaluation on over 9,000 real-world bugs from six diverse Java projects and the GHRB dataset of 131 recent bugs across 16 projects shows that GenLoc consistently outperforms traditional IR-based methods, deep learning models and recent LLM-based techniques in terms of Accuracy@k, MRR@10 and MAP@10. Moreover, GenLoc can localize the highest number of unique bugs, demonstrating its potential to complement existing approaches.
In the future, GenLoc could be extended to support finer-grained bug localization at the method or statement level to provide more precise debugging assistance. Another promising direction is to integrate GenLoc with automated program repair pipelines to evaluate its effectiveness in guiding correct patch generation.